# A spectral method for multi-view subspace learning using the product of projections

### v1

@@ 3.2. Product of projection matrices in a noisy setting
This theorem quantifies the accuracy of estimated subspaces when the ranks are correctly specified. Since the spectral norm of the difference between two projection matrices equals the sine of the largest principle between the corresponding subspaces, our bounds measure the sine of the largest principle angles between $\widehat{\mathcal{J}}$ and $\mathcal{J}$, and between ^$\widehat{\mathcal{I}}_{k}$ and ${\mathcal{I}}_{k}$, respectively. For joint subspace estimation, the bound depends on the perturbation term $R_{\mathcal{J}}$ and the maximum alignment $\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})$ between the individual subspaces. As in Theorem 3.1, large perturbations and subspace alignment can lead to significant errors, even with the correct rank.
3.3. Practical considerations and proposed algorithm
@@
In practice, our primary goal is to use the observed spectrum of $P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$ to accurately determine the joint rank. Our new theoretical insights characterize spectrum perturbation in terms of $\varepsilon_{1}$ and $\varepsilon_{2}$, both of which are unknown in practice. Furthermore, the magnitudes of $\varepsilon_{1}$, $\varepsilon_{2}$ depend on the quality of estimates $\widehat{X}_{k}$. Therefore, we focus on the following practical questions: (i) how can we obtain estimates of $\operatorname*{col}(\widehat{X}_{k})$ for the signal $X_{k}$ in each view $k\in\{1,2\}$ to minimize the magnitude of these perturbations? (ii) how can we estimate $\varepsilon_{1}$ in practice to identify the top cluster of potential joint directions? (iii) how can we bound $\varepsilon_{2}$ to ensure the top cluster contains only joint directions and avoid overlap from random noise alignment? Below, we outline answers to these practical questions and summarize our proposed method.
@@
(i) Estimation of $\operatorname*{col}(X_{k})$. To estimate the signal $X_{k}$, we construct $\widehat{X}_{k}$ using truncated singular value decomposition of observed $Y_{k}$ for each view $k\in\{1,2\}$. We choose the level of truncation based on the criterion in Gavish & Donoho (2014). Let $\beta_{k}=n/p_{k}$ be the aspect ratio, and $y_{\text{med}}$ be the median singular value of $Y_{k}$. Then we only keep those singular vectors whose corresponding singular values exceed the threshold $(0.56\beta_{k}^{3}-0.95\beta_{k}^{2}+1.82\beta_{k}+1.43)\widehat{\sigma}_{k}$. This threshold is designed to obtain an optimal (in mean squared error) reconstruction of the signal matrix. However, it tends to overestimate the true signal rank. As we discuss after Theorem 3.1, overestimating the rank of each view is preferred over underestimation. However, severe overestimation can result in high alignment between random noisy directions, leading to cluster overlap. We will revisit this point in part (iii) and the diagnostic plot generated at the end of the algorithm.
(ii) Estimation of $\varepsilon_{1}$ via rotational bootstrap. From Theorem 3.1, the number of singular values of $\widehat{M}$ above $1-\varepsilon_{1}$ can be used as an estimate of the joint subspace rank. To estimate $\varepsilon_{1}$, we propose a novel bootstrap procedure. We first review a rotational bootstrap approach in Prothero et al. (2024) and then describe our proposal.
Prothero et al. (2024) adopt model (1), and estimate the maximal principle angle between the true subspace $\operatorname*{col}(X_{k})$ and the subspace formed by truncated singular value decomposition of observed $Y_{k}$, $\widehat{X}_{k}=\widehat{U}_{k}\widehat{\Sigma}_{k}\widehat{V}_{k}^{\top}$, $k\in\{1,2\}$, with $\widehat{r}_{k}:=\operatorname*{rank}(\widehat{X}_{k})$. For each bootstrap sample, the data replicate is generated as $Y_{k,(b)}:=U_{k,(b)}\widehat{\Sigma}_{k}V^{\top}_{k,(b)}+\widehat{E}_{k}$, where $\widehat{\Sigma}_{k}$ is a matrix of singular values from $\widehat{X}_{k}$, $U_{k,(b)}\in\mathbb{R}^{n\times\widehat{r}_{k}},V_{k,(b)}\in\mathbb{R}^{p\times\widehat{r}_{k}}$ are orthonormal, drawn uniformly and independently from the Haar measure, and $\widehat{E}_{k}$ is the adjusted estimate of the noise matrix (see Appendix C in Prothero et al. (2024)). Here $X_{k,(b)}=U_{k,(b)}\widehat{\Sigma}_{k}V^{\top}_{k,(b)}$ represents the signal replicate. Using data replicate $Y_{k,(b)}$, the corresponding truncated singular value decomposition of rank $\widehat{r}_{k}$ is given by $\widehat{X}_{k,(b)}:=\widehat{U}_{k,(b)}\widehat{D}_{(b)}\widehat{V}_{k,(b)}^{\top}$. The resulting projection matrices $P_{\operatorname*{col}(X_{k}),(b)}:=U_{k,(b)}U^{\top}_{k,(b)}$ and $P_{\operatorname*{col}(\widehat{X}_{k}),(b)}:=\widehat{U}_{k,(b)}\widehat{U}^{\top}_{k,(b)}$ represent one bootstrap sample of the true and estimated signal subspaces, respectively, allowing estimation of principal angles over replications.
In principle, the above procedure could be applied to estimate $\varepsilon_{1}$. However, it leads to underestimation of the true value, especially when the estimated ranks are small relative to the ambient dimension, as randomly sampled subspaces are likely to be nearly orthogonal; see Appendix B.2 for numerical demonstrations. This underestimation is because the procedure in Prothero et al. (2024) is performed separately for each view $k$. Thus, the alignment across views is not preserved.
The main novelty of the proposed bootstrap procedure lies in preserving the spectrum of the product of projections $P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$ in each bootstrap replicate, in addition to preserving the individual spectrum of $\widehat{X}_{k}$. Specifically, when generating bootstrap samples for each view, we preserve the alignment between the two subspaces rather than sampling uniformly at random. Let $\Sigma_{\widehat{M}}$ be a diagonal matrix containing singular values of $\widehat{M}=P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$. To generate $U_{1,(b)}$ and $U_{2,(b)}$ randomly but with pre-specified alignment, we first generate an $n\times n$ matrix $N$ with entries drawn independently from a standard normal distribution. We let $U_{1,(b)}$ consist of the first $\widehat{r}_{1}$ left singular vectors of $N$ and $U_{2,(b)}$ consist of the next $\widehat{r}_{2}$ left singular vectors. We then align these matrices by updating $U_{2,(b)}$ to be $U_{1,(b)}\cos(\Sigma_{\widehat{M}})+U_{2,(b)}\sin(\Sigma_{\widehat{M}})$. By construction, the singular values of $U_{1,(b)}^{\top}U_{2,(b)}$ are exactly $\cos(\Sigma_{\widehat{M}})$. We then proceed as in Prothero et al. (2024), leading to projection matrices $P_{\operatorname*{col}(X_{k}),(b)}:=U_{k,(b)}U^{\top}_{k,(b)}$ and $P_{\operatorname*{col}(\widehat{X}_{k}),(b)}:=\widehat{U}_{k,(b)}\widehat{U}^{\top}_{k,(b)}$, $k=1,2$, for each replicate. Given $B$ bootstrap replicates, we estimate $\varepsilon_{1}$ as:
@@
where $\Delta_{k,(b)}=P_{\operatorname*{col}(X_{k}),(b)}-P_{\operatorname*{col}(\widehat{X}_{k}),(b)},k\in\{1,2\}$. In Appendix B.3, we show numerically that $\widehat{\varepsilon}_{1}$ tends to overestimate $\varepsilon_{1}$ and becomes more accurate with a larger signal-to-noise ratio.
(iii) Filtering out noise directions via bound on $\varepsilon_{2}$. Due to a low signal-to-noise ratio or rank over-specification, the estimated subspaces $\operatorname*{col}(\widehat{X}_{1}),\operatorname*{col}(\widehat{X}_{2})$ can include many noise directions. These noise directions can exhibit alignment across the two views by random chance, with the degree of alignment increasing as the number of noise directions grows (due to decreasing signal-to-noise ratio and/or increasing rank over-specification). Such random alignments are reflected by the third cluster of singular values in Theorem 3.1. As $\varepsilon_{2}$ becomes large, this noise cluster will overlap with the top cluster corresponding to joint directions, preventing accurate identification of the joint signal. To diagnose such cases in practice, we could adapt a similar bootstrap estimation idea to $\varepsilon_{2}$ as we did to $\varepsilon_{1}$ in (ii). However, we found that a simpler analytical bound can be obtained by utilizing results from random matrix theory.
To motivate the analytical bound, consider a simplified scenario with no individual signals, $P_{\operatorname*{col}(X_{K})}=P_{\mathcal{J}}$, with $P_{\operatorname*{col}(\widehat{X}_{k})}=P_{\mathcal{J}}+P_{\text{noise}_{k}}$, that is the estimated signal is exactly the joint signal plus orthogonal noise directions. Then $\varepsilon_{2}=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}=\left\|P_{\text{noise}_{1}}P_{\text{noise}_{2}}\right\|_{2}$, representing the maximal singular value of the product of random projections. While $\varepsilon_{2}$ is generally more complex, the key intuition remains: only singular values above the threshold expected from random noise alignment can be confidently identified as the signal.
@@
To filter out the noise directions when determining the joint rank, we propose considering only singular values with squares exceeding $\lambda_{+}^{1/2}$. Since, in practice, the true joint rank is unknown, we use marginal $\operatorname*{rank}(\widehat{X}_{k})$ in place of $r_{k}$ to determine the aspect ratios $q_{k}$ in calculating $\lambda_{+}$. This approach can be viewed as conservative, as it effectively assumes that all the directions in estimated $\operatorname*{col}(\widehat{X}_{1}),\operatorname*{col}(\widehat{X}_{2})$ are potentially noisy. Further, we illustrate how these bounds can be used to diagnose model performance in practice.
@@ 4. Simulations on Synthetic Data
Table 1. Mean F-scores (6) on simulated data, larger values correspond to more accurate estimation of joint and individual subspaces. All values are multiplied by $10$. The results are averaged across 50 iterations. The maximum standard error of the mean is $0.03$. We highlight the best result in bold and underline second best result. We denote by () methods that do not allow supplying marginal ranks for each view. We denote Lock et al. (2013) as JIV, Feng et al. (2018) as AJI, Gaynanova & Li (2019) as SLI, Shu et al. (2019) as DCC, Park & Lock (2020) as UNI, and our method as PPD.
@@ 5.1. Colorectal cancer data
Figure 2. Left: Scree plot for RNASeq and mRNAseq data. Blue point denotes the cutoff suggested by Gavish & Donoho (2014). Red point indicates our suggested cutoff. Middle / Right: Spectrum of the product of projection matrices onto each view for two different marginal ranks. The blue region indicates the random direction bound, $[0,\lambda_{+}^{1/2}]$, used to filter out the noise directions. The green region is given by $[1-\widehat{\varepsilon}_{1},1]$ where $\widehat{\varepsilon}_{1}$ is a bootstrap estimate of $\varepsilon_{1}$.
Table 2. Mean cancer subtype misclassification rates in percentages together with the minimum principle angle between estimated individual subspaces. We denote principle components analysis by PCA, Lock et al. (2013) as JIV, Feng et al. (2018) as AJI, Gaynanova & Li (2019) as SLI, Shu et al. (2019) as DCC, Park & Lock (2020) as UNI, and our method as PPD.

### v2

@@ 3.2. Deterministic analysis of the product of projection matrices in a noisy setting
This theorem quantifies the accuracy of estimated subspaces when the ranks are correctly specified. Since the spectral norm of the difference between two projection matrices equals the sine of the largest principle between the corresponding subspaces, our bounds measure the sine of the largest principle angles between $\mathcal{J}$ and $\widehat{\mathcal{J}}$, and between ${\mathcal{I}}_{k}$ and $\widehat{\mathcal{I}}_{k}$ , respectively. For joint subspace estimation, the bound depends on the perturbation term $R_{\mathcal{J}}$ and the maximum alignment $\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}$ between the individual subspaces. As in Theorem 3.1, large alignment between the individual subspaces can lead to significant errors. Note that the error due to such alignment also affects individual subspace estimation through the $\Delta_{\mathcal{J}}$ term in $R_{\mathcal{I}_{k}}$.
3.3. Probabilistic bounds
Theorems 3.1-3.2 are deterministic and hold for any signal estimate $\widehat{X}_{k}$ and for any data-generating noise $Z_{k}$ in (1). In this section, we provide probabilistic bounds for a particular signal estimation procedure and noise model. Specifically, we let $\widehat{X}_{k}$ denote the rank-$\widehat{r}_{k}$ truncated singular value decomposition of $Y_{k}$, and consider Gaussian noise.
Assumption 1.
(Gaussian noise) The entries of $Z_{k}$ in (1) are independently and identically distributed as zero-mean Gaussian random variables with variance $\delta_{k}$, $k=1,2$.
Assumption 2.
(Signal-to-noise ratio) For each $k=1,2$, $\sigma_{r_{k}}(X_{k})\gtrsim\delta_{k}(n^{1/2}+p_{k}^{1/2})$. That is, the smallest non-zero singular value of each view’s signal is sufficiently large.
Under Assumptions 1 and 2, along with rank estimation conditions, we obtain the following probabilistic bounds in terms of the essential aspects of the problem (e.g., the sample size, the number of features in each view, signal ranks) and omit constants. The proof is provided in Appendix A.3.
Theorem 3.3 (Probabilistic bounds).
Suppose Assumptions 1-2 hold. Let $r_{\max}:=\max\{r_{1},r_{2}\}$, $p_{\max}:=\max\{p_{1},p_{2}\}$, $p_{\min}:=\min\{p_{1},p_{2}\}$, $\delta_{\max}:=\max\{\delta_{1},\delta_{2}\}$, and $\sigma_{r}(X):=\min\{\sigma_{r_{1}}(X_{1}),\sigma_{r_{2}}(X_{2})\}$. If $\widehat{r}_{k}\geq r_{k}$ for $k=1,2$, then, with probability greater than $1-\mathcal{O}(\exp\{{-\max(n,p_{\min})}\})$:
$\epsilon_{1}\lesssim\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}.$
If $\widehat{r}_{k}=r_{k}$ then with probability greater than $1-\mathcal{O}(\exp\{-\max(n,p_{\min})\})$:
$\epsilon_{2}\lesssim\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}.$
If additionally $\widehat{r}_{J}=r_{J}$, then with probability greater than $1-\mathcal{O}(\exp\{-\max(n,p_{\min})\})$:
$\max\left\{\left\|\mathcal{P}_{\widehat{\mathcal{J}}}-\mathcal{P}_{{\mathcal{J}}}\right\|_{2},\left\|\mathcal{P}_{\widehat{\mathcal{I}}_{k}}-\mathcal{P}_{{\mathcal{I}}_{k}}\right\|_{2}\right\}$
$\quad\quad\quad\quad\lesssim\frac{1}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}\left\{\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}\right\}.$
Theorem 3.3 makes explicit the distinctive effect of rank estimation on $\varepsilon_{1}$ and $\varepsilon_{2}$. Although we obtained identical bounds, the bound for $\varepsilon_{1}$ holds true for all $\widehat{r}_{k}\geq r_{k}$, whereas the bound for $\varepsilon_{2}$ is only true when $\widehat{r}_{k}=r_{k}$. Intuitively, rank overestimation does not adversely affect $\varepsilon_{1}$, as the alignment of joint directions is already captured. In contrast, adding extra noise dimensions increases the chance of a random strong alignment, pushing $\varepsilon_{2}$ to be high.
Assumption 2 is a signal-to-noise ratio assumption that is common in the random matrix theory due to known concentration $\|Z_{k}\|_{2}\asymp\delta_{k}(n^{1/2}+p_{k}^{1/2})$ (Vershynin, 2025), with high probability. This leads to the standard bound $\|\Delta_{k}\|_{2}\lesssim\delta_{k}(n^{1/2}+p_{k}^{1/2})/\sigma_{r_{k}}(X_{k})$ via Davis–Kahan theorem (Davis & Kahan, 1970). Our bounds on $\varepsilon_{k}$, however, are tighter. Specifically, the first term does not depend on the number of features $p_{k}$, and the dependence on $p_{k}$ is only in the second term, which is of smaller order when the signal $\sigma_{r_{k}}(X_{k})$ is large. Instead, we have the explicit dependence on the rank $r_{k}$, which highlights how smaller-dimensional signals are easier to estimate.
The third bound highlights the difficulties that arise from the non-orthogonal individual subspaces $\mathcal{N}_{k}$. As expected, high alignments across these subspaces makes the separation of joint and individual structures more difficult, affecting estimation accuracy.
4. Proposed Algorithm Based on Product of Projection Matrices
@@
In practice, our primary goal is to use the observed spectrum of $P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$ to accurately determine the joint rank. Our new theoretical insights characterize spectrum perturbation in terms of $\varepsilon_{1}$ and $\varepsilon_{2}$, both of which are unknown in practice and depend on the estimates $\widehat{X}_{k}$. Specifically, Theorem 3.1 tells that singular values above $1-\varepsilon_{1}$ correspond to the joint, while those below $\varepsilon_{2}$ correspond to the noise directions. Therefore, we focus on the following practical questions: (i)  how to estimate $\operatorname*{col}(\widehat{X}_{k})$ for the signal $X_{k}$ in each view $k\in\{1,2\}$? (ii) how to estimate $\varepsilon_{1}$ in practice to identify the top cluster of potential joint directions? (iii) how to bound $\varepsilon_{2}$ to ensure the top cluster contains only joint directions and avoid overlap from random noise alignment? From the discussion in Section 3.2, we also have the following insights: (i) rank overestimation is preferred to underestimation since it leads to smaller $\varepsilon_{1}$; (ii) simultaneously, overestimation pushes $\varepsilon_{2}$ closer to 1 which makes it harder to distinguish the joint signal from the noise; (iii) rank-to-dimension ratio, signal-to-noise (snr) ratio, and individual component alignment all have influence on the final estimation quality – in general, high signal rank, low snr, and high degree of alignment make the problem highly intractable. Below, we outline answers to these practical questions and summarize our proposed method.
@@
(i) Estimation of $\operatorname*{col}(X_{k})$. To estimate the signal $X_{k}$, we construct $\widehat{X}_{k}$ using truncated singular value decomposition of observed $Y_{k}$ for each view $k\in\{1,2\}$, an approach which allowed us to derive probabilistic bounds in Section 3.3. Since overestimation is preferred over underestimation, we choose the level of truncation based on the criterion in Gavish & Donoho (2014), which tends to overestimate the signal rank. Let $\beta_{k}=n/p_{k}$ be the aspect ratio, and $y_{\text{med}}$ be the median singular value of $Y_{k}$. Then we only keep those singular vectors whose corresponding singular values exceed the threshold $(0.56\beta_{k}^{3}-0.95\beta_{k}^{2}+1.82\beta_{k}+1.43)\widehat{\sigma}_{k}$. This threshold is designed to obtain an optimal (in mean squared error) reconstruction of the signal matrix. Because severe overestimation is still problematic as it introduces potentially many noise directions, we show how to filter out these directions in part (iii) while also we provide a diagnostic plot that allows to confirm the choice.
(ii) Estimation of $\varepsilon_{1}$ via rotational bootstrap. From Theorem 3.1, the number of singular values of $\widehat{M}=P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$ exceeding $1-\varepsilon_{1}$ provides an estimate for the joint rank, $r_{\mathcal{J}}$. This approach, however, requires an estimate for the threshold parameter $\varepsilon_{1}$. To this end, we propose a novel rotational bootstrap procedure. Previous bootstrap procedures, e.g., in Prothero et al. (2024), independently resample each view’s signal subspace, $U_{k}$. This fails to preserve the alignment between $U_{1}$ and $U_{2}$, since in high-dimensional settings, independently sampled subspaces are likely to be nearly orthogonal, particularly when the rank-to-dimension ratio is small. As a result, independent bootstraping leads to underestimation of $\varepsilon_{1}$, and consequently, an unreliable estimate of the joint rank.
[…]
@@
where $\Delta_{k,(b)}=P_{\operatorname*{col}(X_{k}),(b)}-P_{\operatorname*{col}(\widehat{X}_{k}),(b)},k\in\{1,2\}$.
(iii) Filtering out noise directions via bound on $\varepsilon_{2}$. When the estimated subspaces $\operatorname*{col}(\widehat{X}_{1})$ and $\operatorname*{col}(\widehat{X}_{2})$ contain noise directions—a consequence of low signal-to-noise ratio or rank over-specification—these directions can exhibit spurious alignment purely by chance. Such random alignments produce a cluster of singular values bounded from above by $\varepsilon_{2}$, as established in Theorem 3.1. A large value of $\varepsilon_{2}$ complicates the separation of joint and noise subspaces, as it implies a potential overlap in their corresponding singular value distributions.
To filter out noise directions, one could adapt the bootstrap methodology to estimate $\varepsilon_{2}$ and use it as a filtering threshold. However, we find a simpler and more direct approach to be effective. By leveraging results from random matrix theory, we found a deterministic upper bound for $\varepsilon_{2}$ that depends only on the rank-to-dimension ratios.
To motivate the analytical bound, consider a simplified scenario with no individual signals, $P_{\operatorname*{col}(X_{K})}=P_{\mathcal{J}}$, with $P_{\operatorname*{col}(\widehat{X}_{k})}=P_{\mathcal{J}}+P_{\text{noise}_{k}}$, that is the estimated signal is exactly the joint signal plus orthogonal noise directions. Formally, $P_{\text{noise}_{k}}$ is the projection onto the subspace $\mathcal{J}^{\perp}\cap\operatorname*{col}(\widehat{X}_{k})$. Then $\varepsilon_{2}=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}=\left\|P_{\text{noise}_{1}}P_{\text{noise}_{2}}\right\|_{2}$, representing the maximal singular value of the product of random projections. While $\varepsilon_{2}$ is generally more complex, the key intuition remains: only singular values above the threshold expected from random noise alignment can be confidently identified as the signal.
@@
This analytical result provides a practical filtering strategy. We propose to classify only those singular values as joint signal whose magnitude exceeds the spectral edge $\lambda_{+}$ (in addition to exceeding $1-\widehat{\varepsilon}_{1}$). In practice, we calculate the aspect ratios $q_{k}$ using the full estimated marginal ranks, $\operatorname*{rank}(\widehat{X}_{k})$. This substitution makes the bound conservative, as it provides a worst-case noise estimate by assuming all non-joint directions could be noise.
@@ 5. Simulations on Synthetic Data
Table 1. Mean F-scores (6) on simulated data, larger values correspond to more accurate estimation of joint and individual subspaces.
@@ 6.1. Colorectal cancer data
Figure 2. Left: Scree plot for RNASeq and mRNAseq data. Blue point denotes the cutoff suggested by Gavish & Donoho (2014). Red point indicates our suggested cutoff. Middle / Right: Spectrum of the product of projection matrices onto each view for two different marginal ranks. The blue region indicates the random direction bound, $0,\lambda_{+}^{1/2}]$, used to filter out the noise directions. The green region is given by $1-\widehat{\varepsilon}_{1},1]$ where $\widehat{\varepsilon}_{1}$ is a bootstrap estimate of $\varepsilon_{1}$.
Table 2. Mean cancer subtype misclassification rates in percentages together with the minimum principle angle between estimated individual subspaces.